OpenStudy (anonymous):

WILL FAN AND GIVE METAL: Exactly 5.00 L of air at -50.0*C is warmed to 100.0*C. What is the new volume if the pressure remains constant?

OpenStudy (somy):

V1 / T1 = V2 / T2 V2 is unknown so find it, the rest is given by the question but, convert the Celsius to Kelvin by adding 273 to Celsius value

OpenStudy (anonymous):

so then I would get 223 and 273? @Somy

OpenStudy (somy):

-50 to Kelvin, yes 223

OpenStudy (somy):

100 to kelvin would be 273+ 100 = K